OpenStudy (anonymous):

Radical expressions? I REALLY NEED HELP. I DON'T UNDERSTAND MATH D:

OpenStudy (anonymous):

\[\sqrt{9y+27}+\sqrt{4y+12}\]

OpenStudy (anonymous):

\[11y \sqrt{3}\]?

OpenStudy (anonymous):

*13y

OpenStudy (dan815):

im here to rescue

OpenStudy (anonymous):

omgg ty ty. what did i do wrong??

OpenStudy (dan815):

you cannot add whats in different sqrts like that

OpenStudy (dan815):

lets just simplfy first

OpenStudy (anonymous):

\[9\sqrt{3y}+4\sqrt{3y}\]

OpenStudy (dan815):